OpenStudy (anonymous):

Convert the equation to the standard form for a parabola by completing the square on x or y as appropriate. y2 + 2y - 2x - 3 = 0 A. (y + 1)2 = 2(x + 2) B. (y - 1)2 = -2(x + 2) C. (y + 1)2 = 2(x - 2) D. (y - 1)2 = 2(x + 2)

OpenStudy (anonymous):

@cobra-strikes

OpenStudy (dumbcow):

when completing the square, take half of middle coefficient, square it, then add to both sides example \[y^2 + 8y = 0\] \[y^2 +8y + (\frac{8}{2})^2 = (\frac{8}{2})^2\] \[y^2 +8y +16 = 16\] \[(y+4)^2 = 16\]